OpenStudy (anonymous):

How many moles of H3PO4 are produced when 20.0g of HCl are produced by the reaction PCl5 + 4H2O -> H3PO4 + 5HCl? again Im looking for the mole

OpenStudy (chmvijay):

hey first convert 20 gram into moles then i will tell u u know the equation

OpenStudy (chmvijay):

did u get me

OpenStudy (chmvijay):

if 5 moles of HCl produces - 1 mole of phosphoric acid then x moles of Hcl produces =(1*x)/5

OpenStudy (chmvijay):

X=20/36.8=0.5434 moles of HCL =0.5434*1/5 =0.108 moles of phosphoric acid will be produced

OpenStudy (anonymous):

well, that was actually the answer i had. but I guess the answer my instructor wanted was the set up of the equation.

OpenStudy (anonymous):

Also, my calculation was a little off.

OpenStudy (anonymous):

0.09 moles.

OpenStudy (chmvijay):

ok ok do u need any clarification

OpenStudy (anonymous):

no, that is all. Thanks for your help.

OpenStudy (chmvijay):

yaaa 0.09 moles nearly euals to 0.1 i do agrree with u muhammad